### v1

# Improved Contact Graph Routing in Delay Tolerant Networks with Capacity and Buffer Constraints

Abstract
Satellite communications present challenging characteristics. Continuous end-to-end connectivity may not be available due to the large distances between satellites. Moreover, resources such as link capacity and buffer memory may be limited. Routing in satellite networks is therefore both complex and crucial to avoid packet losses and long delays. The Delay Tolerant Network (DTN) paradigm has emerged as an efficient solution for managing these challenging networks. Contact Graph Routing (CGR), a deterministic routing algorithm, is one of the most popular DTN algorithms. CGR is compatible with the “store, carry, and forward” principle, whereby a node receives a message and stores it in its buffer until a transmission opportunity becomes available. However, CGR relies on simplified models to incorporate potential constraints in the route search. For instance, the linear volume assumption is often used to consider capacity constraints. Moreover, capacity management and buffer management are mostly performed during the forwarding phase, once an issue has occurred. In this paper, we propose to take measures before or during the route search in order to find routes that respect both contact-capacity limits and node-buffer limits. We introduce the contact splitting and edge pruning operations to effectively account for the routing constraints. This ensures that CGR outputs the optimal solution among the subset of valid solutions. The proposed approach can also be used to book resources to be used in case of issues during the forwarding step.
Index Terms:
Delay tolerant networks, contact graph routing, satellite communications, buffer, contact capacity.
I Introduction
Satellite communications belong to the category of Intermittently Connected Networks (ICNs). They are characterized by intermittent connectivity between the nodes of the network, preventing continuous end-to-end connections and resulting in high delays. This intermittent nature induces multiple discontinuous episodes of communication opportunities between pairs of nodes (so called contacts, see the following sections). A transmission can thus be delayed if there is no immediate communication opportunity. Nodes must therefore operate on a “store, carry, and forward” principle.
As a result, since traditional Internet protocols are not suited to those challenges, the need for new protocols arose. The Delay Tolerant Network (DTN) [1] architecture with a new protocol layer has been proposed by the Internet Engineering Task Force (IETF) to handle these issues. The Bundle Protocol (BP) [2], considered by the Consultative Committee for Space Data Systems (CCSDS), is an instance of such protocols.
As mentioned above, routing is challenging with such discontinuous and time-dependent connections. The main routing issue in ICNs is the inability to establish a complete route due to lack of connectivity. Moreover, in addition to the delivery-time objective, optimizing delivery rate and resource consumption should also be considered. For instance, the use of transmission resources must be carefully optimized because of their limited availability. Additionally, as data units are often held in the node buffers for long periods, optimizing storage resources to avoid congestion and loss is crucial. Efficient routing algorithms are therefore needed in addition to suitable network capabilities via the above-mentioned protocols.
Routing algorithms in DTNs can be classified into three categories depending on the availability of network information. They can be opportunistic [3], probabilistic [4] or deterministic, depending on whether the contacts are unpredictable, determined with a certain probability, or completely predetermined, respectively. In this paper, we focus on space communications, where natural and artificial satellites have deterministic motions. Thus, contacts can be predetermined, making deterministic routing algorithms suitable. The widely used deterministic routing algorithm in DTNs is the Contact Graph Routing (CGR) [5].
Since its invention, CGR has gained attention in the research community and several improvements have been proposed. For instance, using Dijkstra’s algorithm for the route search and choosing the Best-case Delivery Time (BDT) as objective were presented in [6]. Then, the Earliest Transmission Opportunity (ETO) was introduced to compute the time at which a bundle is transmitted through the contact [7]. The ETO takes into account the time a bundle has to wait before being transmitted because of other scheduled bundle transmissions on the same contact. Later, a complementary overbooking management was added to the ETO to improve routing decisions [8]. These enhancements among others result in the currently used version of CGR [5]. Nevertheless, this version provides only local and reactive contact-capacity and buffer-occupancy managements, which leads to sub-optimal solutions. Therefore, it is important to explore alternative approaches.
One modern application, where capacity and buffer managements may be critical, is cislunar communications. With the advent of the Artemis Program [9], many projects in the lunar environment arose, triggering a high need of communication resources [10, 11]. Another application is low-earth orbit satellite constellations: These constellations are expected to be utilized extensively in the future. However, not all future constellations will be large enough to maintain end-to-end connectivity at all times. Even large constellations may not deploy all their satellites at the initial stage, requiring the establishment of a network with a limited number of satellites during the initial phase. As a result, many constellations are expected to temporarily buffer data.
Main contributions. In this paper, we provide a routing solution based on CGR with a global and proactive congestion control. We propose to perform contact-capacity management and buffer-limit management before or during the route search such that the obtained route is optimal and respects these two constraints. The objective is achieved as follows.
  • We propose managing the contact capacity before the route search by modifying the network information, such as the provided contact lists, using contact splitting. Contact splitting removes the part of the contacts used by a routed bundle.
  • We propose to track node-buffer occupancy as function of time using forecast buffer tables. The buffer occupancy is then managed by:
    • Temporarily modifying the network information, also via contact splitting, before routing a specific bundle to avoid buffer overflows.
    • Temporarily forbidding some edges between contacts during the route search for a specific bundle to avoid buffer overflows.
The modified network information enables to remove the routes that do not respect the constraints (and only these routes). Subsequently, the route-search algorithm, such as CGR, can find the optimal solution among the subset of valid solutions.
The proposed approach can also be considered to book resources to be utilized in case of failure events during the forwarding phase. These resources, referred to as safety margin in the paper, are to be used by intermediate nodes having limited information on the network state.
Structure of the paper. The paper is structured as follows. In Section II, we first introduce key terms used in the paper and we provide a review of CGR. In Section III, we focus on existing capacity an buffer management challenges in DTNs to introduce our motivations. The proposed solution is introduced in Section IV. The benchmark source-routing CGR, considered as benchmark, is detailed in Section V. Simulation assumptions and results comparing our method with the benchmark are given in Sections VI and VII, respectively.
II Background and motivations
II-A Key notions
We first define some key terms used in this paper.
  • Node: In a satellite communication scenario, a node $X$ can be a satellite or a Ground Station (GS) as illustrated in Figure 1.
  • Contact: A contact $C_{j}$ is defined by the communication opportunity between two nodes. It is thus characterized by sender and receiver nodes, a start time $t_{1,j}$, an end time $t_{2,j}$, and a data rate/capacity $R_{C_{j}}$ in bits per second (bps). The volume of a contact is its rate multiplied by its duration.
  • Contact Plan (CP): A CP is the list of contacts for a time period. It reflects the evolution of the network topology over time. An example is provided in Figure 2.
  • Edge: An edge exists between two contacts when the receiver of one contact matches the sender of a second contact, and the end time of the latter is later than the start time of the former.
  • Bundle: A bundle is the data unit encoded by the BP. A bundle $i$ is characterized by its size $S_{i}$ in bits. A bundle is generated at a source node and is provided with a destination node.
  • Route: A route or a path is a series of contacts that are used to deliver the bundle from the source node to the destination node.
Figure 1: Snapshot of a satellite communication scenario at time $t$ where $t_{1,1}\leq t\leq t_{2,1}$ and $t_{1,4}\leq t\leq t_{2,4}$ (see Figure 2).
II-B Contact graph routing
CGR is divided into three main phases [12, 5]. First, the CP, for example constructed at a mission control unit, is forwarded to all nodes in the network. Second, in the forwarding (i.e., transmission) phase, each node receiving a bundle performs the route search using Dijkstra’s algorithm to determine the next hop. Third, checks are performed at each intermediate node before transmission to verify that system constraints are respected.
II-B1 CP construction
In a deterministic network case, the CP is constructed at the source nodes using the network information and transmitted to all intermediate nodes for the route-search phase.
II-B2 Route search
For a given destination, a route between the source node and the destination node is computed using Dijkstra’s algorithm [13]. However, an adaptation of this algorithm is used to consider the discontinuities of ICNs. The main idea is to use a contact-graph representation of the network rather than a node-graph representation. Dijkstra’s algorithm is thus implemented on the contact-graph representation of the network. The output of a Dijkstra search is the route from the source to the destination having the BDT.
Yen’s algorithm [14] enables to determine a list of $k$ shortest routes by iteratively exploring the routes between a source and a destination, using Dijkstra with the contact-graph representation while avoiding loops and previously discovered paths.
In the standard CGR implementation, each node computes a list of $k$ shortest routes to each node in the network using Yen’s algorithm, where $k$ is an input parameter. Then, once a bundle reaches a given intermediate node, the node determines a feasible precomputed route from its route list to determine the next hop.
II-B3 Validation and forwarding
During the forwarding phase, the local queuing state of the buffer and the bundle properties are available. The remaining contact volume and the buffer queuing time for bundles scheduled on the same contact can then be taken into account. These checks, among others, help verify the validity of previously computed routes in the route list and prevent local congestion.
A list of candidate valid routes can thus be determined from the $k$ shortest routes. A valid route should guarantee contact volume availability given the size of the bundle. Moreover, one should also check that, taking into account transmission and queuing time, the bundle is effectively transmitted before the end of the corresponding contact, see ETO [7].
The best route among valid routes is selected to determine the bundle’s next hop. Once a route has been selected, the bundle is queued for transmission, and the available volume of the contact used is decreased by the size of the bundle.
II-B4 Additional comments on CGR
Although widely used, the CGR is computationally complex. Each node has to perform many checks for each bundle to only determine the next contact to use. Moreover, since these checks are performed after the route search, valid solutions may not be found among the candidate routes. This may trigger a time-consuming iterative search process between the Dijkstra algorithm and the checks.
II-C Source-routing CGR
One solution to reduce CGR’s complexity is Source Routing (SR) [15]. Unlike per hop routing, the idea of SR is to compute the route at the source node and encode it into the bundle in an extension block [16].
When the bundle, whose route has already been calculated, arrives at an intermediate node, this node only checks the validity of the encoded route. In case of invalidity, traditional CGR is performed from this intermediate node. The following two types of SR can be implemented.
II-C1 Bundle-independent SR
With bundle-independent SR, a source performs the route search without bundle properties and queuing information. During the forwarding time, the checks of CGR such as ETO and volume availability are performed. The validity of the encapsulated route is checked using available information on bundles, like its size, and local queue states [17].
II-C2 Bundle-dependent SR
When available at the initial route search time, the bundle properties, e.g., size and priority, can be utilized [5], as is the case with bundle-dependent SR. Then, during the forwarding phase, information on local queues is taken into account and the viability of the route is checked.
Our solution goes further in this paradigm by incorporating network information alongside bundle properties during the initial route search.
III Contact-capacity and node-buffer challenges in DTNs
III-A Contact-capacity management
As mentioned previously, in CGR capacity consumption and availability are managed during the forwarding phase. The common approach to take contact capacity into account in the route search is to check that the available volume of contacts is greater than the size of the bundle. This is the so called linear approach [5]. However, this does not prevent exceeding capacity limits: It does not track the time at which the bundle uses the contact. The linear volume model is represented on the right side of Figure 3, see also [5].
We detail this issue with the example of Figure 3. In this example, when using the linear model, one may think that bundle $i$ (hatched rectangle) reaching node C at $t_{i}$ can use contact $C_{3}$ immediately. Nevertheless, in reality the contact volume is consumed as represented on the left side of Figure 3. In this case, a bundle reaching node C at $t_{i}$ cannot be transmitted using contact $\text{C}_{3}$ immediately because bundle $a$ (striped rectangle) is being transmitted and uses all the capacity at this specific time. In traditional CGR, this problem is detected during the forwarding step, after the route search, using the ETO for example. The transmission of the bundle should thus be delayed accordingly.
Nevertheless, delaying the transmission of bundle $i$ in the forwarding phase may cause multiple problems such as inaccurate estimates of delivery time, the end or unavailability of current or future contacts on the planned route, or reaching the bundle delivery deadline before its delivery.
These collision problems can be better managed, or even avoided, if capacity consumption is modeled differently. In the same example, if the use of $\text{C}_{3}$ at that precise moment had been known in advance, an alternative route might have been a more effective solution for delivering bundle $i$.
Figure 2: Initial CP.
Figure 3: Linear vs real volume consumption.
III-B Buffer management
In DTNs, node buffer management is an important task, as bundles can wait a long time in node buffers before being transmitted. In addition, available buffer capacities may be limited.
Buffer management strategies usually consist of addressing a problem once it has occurred. For example when a buffer is full, the decision to reroute or drop bundles takes place on the spot, like in reactive custody transfer [2]. Authors in [18] suggested a strategy to refuse custody in order to preserve buffer space at nodes. However, these strategies rely only on information coming from local neighbor nodes.
An alternative approach would be to have a global vision on buffer occupancy and to handle the buffer space issues during the initial route search. Note however that, unlike the volume and capacity that are managed at the contact level, the buffer occupation is managed at the node level. For instance, multiple bundles using distinct contacts can affect the buffer of the same node. Therefore, one should be careful when addressing this challenge in a contact-graph representation of the network.
IV Proposed global capacity and buffer management solution
The framework considered is the bundle-dependent SR-CGR [15], see Section II-C2. We propose to address the buffer and capacity constraints before or during the route search. As a reminder, this is to be opposed with approaches reported in the state-of-the-art where actions are taken once an issue has occurred.
The solution ensures that any route found respects the buffer and capacity limits. Moreover, the proposed modifications do not remove any viable solution, which enables the route search to find the optimal solution given the constraints.
This is achieved by modifying the CP before starting the route search, both for the capacity and buffer constraints. Moreover, for the buffer constraints, the allowed edges between contacts are also modified during a route search. In other words, we modify the parameters such that CGR performs the search only among the valid solutions.
The optimal route that respects buffer and capacity limits is then encoded in the bundle, which is forwarded through the network. This manner of sharing the bundle transmission path is the same as in source-routing CGR.
IV-A Model-parameter modifications
Our solution consists of two key elements:
  • First, we propose to modify the CP to be used by a route-search algorithm (e.g., Dijkstra search) via contact splitting. We differentiate between two kinds of CP modifications: The definitive ones, which are to be shared in the network, and the temporary bundle-specific modifications, which are performed locally by one node within its route-search process for a specific bundle.
  • Second, we propose imposing additional restrictions on allowed edges between contacts. These rules are then to be respected inside the route-search algorithm. These restrictions are to be implemented only locally by a node, during its route search for a specific bundle.
IV-A1 Contact splitting to modify the CP
We assume that the CP is shared by all entities performing a route search (e.g., source nodes).
We introduce the contact splitting operation. Let us consider an original arbitrary contact $\text{C}_{j}$, having a data rate $R_{C_{j}}$, between a sender node $M$ and a receiver node $N$ from $t_{1,j}$ to $t_{2,j}$. Contact splitting consists in replacing the original contact in the CP, by two split contacts $\text{C}_{j,1}$, $\text{C}_{j,2}$ between nodes $M$ and $N$: one from $t_{1,j}$ to $t_{e_{1}}$ and the other one from $t_{e_{2}}$ to $t_{2,j}$, such that $t_{1,j}\leq t_{e_{1}}\leq t_{e_{2}}\leq t_{2,j}$, see Figure 4. Split contacts have the same data rate $R_{C_{j}}$ as the original contact $C_{j}$.
Note that if $t_{e_{2}}-t_{e_{1}}>0$, part of the original contact is erased. Moreover, one of the split contacts might last for no time, meaning that contact splitting results in contact shortening.
Figure 4: Contact splitting.
IV-A2 Edge pruning
We define edge pruning as the act of forbidding edges between contacts. This action can be taken between any pair of contacts (original or split). For example, in Figure 3, one can choose to forbid the edge between contacts $\text{C}_{1}$ and $\text{C}_{2}$. This means that, during a route search, there is no option to choose a route having contact $\text{C}_{1}$ followed by contact $\text{C}_{2}$.
IV-B Capacity and buffer managements
We explain how contact splitting and edge pruning are used to address the capacity and buffer constraints discussed in Section III.
IV-B1 Capacity management
We assume knowledge of the bundle size. First, the route search for the first transmitted bundle $i$ is performed. A contact is skipped if the bundle size is greater than the contact volume (similarly to the linear-volume approach). Once the shortest path is selected for this bundle, the CP is modified as follows: We remove the part of the contact consumed by the bundle from the CP before performing the route search for subsequent bundles. The length of the contact to be erased is $t_{\text{Tx,}i}^{C_{j}}$, starting from the start time of the contact or the arrival time of the bundle at the contact’s sender node, whichever occurs later. This variable $t_{\text{Tx,}i}^{C_{j}}$ represents the time needed to transmit bundle $i$ over contact $C_{j}$ in the selected path:
$t_{\text{Tx,}i}^{C_{j}}=\frac{S_{i}}{R_{C_{j}}}.$ (1)
This operation is a contact splitting operation as it creates two new contacts. These two new contacts are to be added to the CP and the original contact is to be removed. This operation is performed for all contacts in the selected route. Then, the subsequent route searches use this modified CP.
The CP modification, to be performed after each selection of a route for a bundle, is definitive and has to be shared with all nodes performing route searches.
Example. In the example of Figure 3, after route search and route selection for the initial bundle $a$ (striped rectangle), the CP becomes as represented in Figure 5. The new CP is then used for route search of subsequent bundles. In this case, the part of contact $\text{C}_{3}$ between $t_{i}$ and $t_{i}+t_{\text{Tx,a}}^{C_{3}}$ is not an option for future bundle and no collision will therefore occur.
Figure 5: Modified CP after allocating route for bundle $a$.
IV-B2 Buffer management
We first introduce the notion of forecast buffer tables. The goal of these tables is to store the buffer occupancy status of each node as a function of time, including future time. Similarly to the CP, we assume that these tables are shared by all entities performing route searches. We then explain how to use these tables to respect the buffer constraints.
Forecast buffer tables. We assume that each node has a maximum buffer capacity limit $B_{\text{max}}$ and that an overflow occurs if this limit is exceeded.
The proposed buffer management solution uses temporary bundle-specific CP modifications, not shared in the network (see the following subsection). However, once a route is selected, the forecast buffer tables are updated accordingly and shared to all entities performing route search. In other words, once a route is validated, one checks when the bundle will reach a given node and how long it will stay in its buffer according to the validated route.
Since the buffer tables are updated as soon as a route is validated, all routing nodes know in advance the status of the buffers as function of time, even before the bundles travel through the network.
Algorithm to respect the buffer constrains. The main ideas of the proposed algorithm to enforce buffer-limit constraints are as follows. Since the bundle size and the forecast buffer tables are available at the route-search time, one can check whether the bundle can be in the buffer of any node at any time. If not, the algorithm must forbid routes leading to the storage of the bundle in potentially problematic buffers. Hence, before performing a route search, the CP is temporarily modified using contact splitting. Then, some edges are pruned using edge pruning inside the route-search algorithm. This algorithm can be formulated with the four following main steps.
Step 1: To begin with, the size of the bundle is virtually added to all node buffer tables at all times. Potential overflows are detected and determined by their location (which nodes) and time (start and end times). The set of nodes $Z_{k}$ with potential overflows for bundle b is denoted by $\mathcal{Z}_{\text{b}}$. Each $Z_{k}\in\mathcal{Z}_{\text{b}}$ has a number of $n_{\text{of}}\geq 1$ potential overflows forming a set of overflows denoted by $\mathcal{O}_{Z_{k}}$. The start and end times of each overflow are denoted by $t_{1,o_{l}}$ and $t_{2,o_{l}}$, $o_{l}\in\mathcal{O}_{Z_{k}}$, respectively.
Example. Figure 6 presents an example where a bundle b arrives at a source node and needs to be routed. The figure depicts one node $Z_{1}\in\mathcal{Z}_{\text{b}}$ having two potential overflows $o_{1}\text{ and }o_{2}\in\mathcal{O}_{Z_{1}}$. These potential overflows are identified by virtually adding the size of bundle b (vertically striped rectangles) to the buffer level of node $Z_{1}$ (dotted rectangles). The buffer level of node $Z_{1}$, before virtually adding the size of bundle b, is determined by validated routes of previous bundles (dotted rectangles). The corresponding forecast buffer table for node $Z_{1}$ is shown in Figure 7.
Figure 6: Potential overflows at node $Z_{1}$.
Figure 7: Forecast buffer table of node $Z_{1}$ before routing bundle b.
Figure 8: Contacts which might result in overflow $o_{1}$ at node $Z_{1}$.
Step 2: Then, the original CP is temporarily modified accordingly (before starting the route search). One should remove all parts of the contacts from the CP that might forward bundle b to any node $Z_{k}\in\mathcal{Z}_{\text{b}}$ between overflows start and end times using contact splitting. These problematic contacts are contacts that start before the end of an overflow and end after its start.
Example. Some problematic contacts for the potential overflow $o_{1}$ descried in the previous example are shown in Figure 8. These contacts have node $Z_{1}$ as receiver (regardless of their sender node $X$), start before $t_{2,o_{1}}$, and end after $t_{1,o_{1}}$. They are problematic because their duration intersects the period of overflow $o_{1}$. The problematic portions of these contacts are horizontally striped in Figure 8. They may forward bundle b to $Z_{1}$ during the potential overflow $o_{1}$ duration. The parts of these contacts that intersect the overflow duration are to be temporarily removed using contact splitting. The same applies to all potential overflows at all nodes.
Step 3: Removing the problematic contact’s potential overflow parts is not sufficient. One should also address problematic contact successions. These problematic contact successions occur between contacts having an end time before the beginning of the potential overflow $o_{i}$ and a succeeding contact in a route having a start time after the beginning of overflow $o_{i}$. Such succession of contacts is problematic because it results in having the bundle in the buffer of the node during the potential overflow. This step therefore consists in identifying the problematic contact successions as described via the following example.
Example. After step 2, there is no contact having $Z_{1}$ as receiver that spans the duration of $o_{1}$ in Figure 8. Nevertheless, there are contacts having $Z_{1}$ as receiver and ending before $t_{1,o_{1}}$. Using one of these contacts (e.g., $C_{i}$), a bundle arrives to $Z_{1}$ before the potential overflow $o_{1}$. However, this bundle must leave $Z_{1}$ before the beginning of overflow $o_{1}$ to avoid the potential problem. The bundle leaves node $Z_{1}$ through a contact succeeding the previous one ($C_{i}$) in a route. The succeeding contact (e.g., $C_{j}$) has $Z_{1}$ as sender, regardless of the receiver. As a result, the succeeding contact $C_{j}$ should start before overflow $o_{1}$ to avoid it. Otherwise, the succession of $C_{i}$ and $C_{j}$ in a route causes the overflow and is therefore problematic.
Step 4: Finally, the last step is to forbid edges between contacts yielding problematic contact successions. In other words, the edge with the contacts starting too late are pruned. This edge pruning is performed during the route search.
Example. In the example of Figure 8, during the route search, if a contact $C_{i}$ forwards the bundle to node $Z_{1}$ before $o_{1}$, the edge between $C_{i}$ and a contact $C_{j}$ from $Z_{1}$ to a third node is only accepted if $C_{j}$ starts before $t_{1,o_{1}}$. The same applies to all overflows at all nodes.
As described by the above steps, only invalid routes are removed by the modifications of the proposed contact capacity and buffer management solutions. In other words, the proposed modifications before and during the route search do not omit any feasible solution. Hence, the subsequent use of CGR gives the optimal route in the subset of valid solutions.
IV-C Booking resources for forwarding safety margin
With the proposed solution only source nodes keep track of the buffer and contact capacity occupancy or status. Due to this feature, they can avoid overflows and overbooking. Nevertheless, due to some unpredictable disruptions, like antenna pointing errors or node power outages [19], intermediate nodes may have to perform routing during forwarding time. Since intermediate nodes may not have access to buffer and capacity status, their routing decision may cause overflows or overbooking. For this reason, we propose to let parts of the original contact and buffer resources unused by source nodes. These unused portions are referred to as safety margins and are utilized by intermediate nodes to reroute bundles without resulting in overflows or overbooking. The quantity of resources to be allocated through this method can be determined as a function of the network’s disruption level.
These safety margins can be implemented as follows. First, regarding the contact-capacity management, two CPs are created. Contact splitting, for the slots that should be kept as margin, is performed. The safety margin is therefore removed from the original CP. The obtained CP is provided to and used by the source nodes. The safety margin part (contacts removed from the original CP) is forwarded to intermediate nodes that may perform routing in unpredictable situations.
In terms of buffer management, source nodes utilize a fraction of the buffer capacity, leaving the rest available for intermediate nodes. This solution therefore enhances robustness to potential unexpected events, despite the poor knowledge of the current network status by intermediate nodes.
IV-D Information sharing in the network
As mentioned above, the modified network information after contact splitting (except the bundle-specific CP modifications) and the modified forecast buffer tables (after allocating a route for a bundle) are to be shared to all source nodes performing a route search. This is similar to the sharing of the up-to-date CP to the source nodes in standard SR-CGR.
The safety margin resources are shared to all nodes in the network at a low frequency.
V Benchmark: source-routing CGR
As mentioned in the first part of the manuscript, network issues are handled after the route search in standard SR-CGR. While retransmission mechanism do exist for long round-trip times communication, such as Licklider Transmission Protocol (LTP) [20], retransmission mechanism in case of failure in standard SR-CGR is not well described in the literature. As there are many manners to address these issues, we describe in details the assumptions made for the considered SR-CGR benchmark. The proposed algorithm is compared to this benchmark in Section VII, where the simulation results are presented.
To begin with, both in the benchmark and in our algorithm, when a bundle is being transmitted over a contact, it is assumed that all the link is occupied for the time required to transmit the bundle. In other words, the whole link capacity is used (i.e., Time Division Multiplexing (TDM) approach). The initial CP is constructed at the source and is broadcasted to all nodes in the network.
Then, the source node performs route search using Yen’s algorithm with Dijkstra searches. A list of the $k$ (Yen’s constant) shortest routes is generated. This list is kept until a change occurs in the CP, such as the end of a contact. In this case, the route list is discarded and a new list is generated, as in route list pruning [12]. A list of candidate routes is constructed as explained in Section II-B3. When no candidate route is found in this list, the bundle is delayed and the route search is performed again after one time step.
Regarding the capacity and buffer managements performed before the route search: When a route is found, the volumes occupied by the bundle in the selected route are deducted from the available volume of corresponding contacts in the CP (i.e., linear volume approach as described in Section III-A). No buffer information is available or considered at the source.
Finally, the route is encoded in the bundle and the bundle is sent in the network. A bundle with an encapsulated route at a transmitting node (source or intermediate node) is subject to checks to verify whether this route can be used or not. The contact capacity and buffer managements during the forwarding phase in the benchmark are detailed hereafter.
V-1 Contact-capacity management after the route search
First, the time at which the corresponding contact must be used for the bundle transmission is verified. In case of collision with another bundle, the transmission of the bundle having the highest ID is delayed (similar behavior to that of determining the ETO). The ID increases with the bundle generation time at the source. Two cases can occur:
  • If delaying the bundle does not cause any problem, the times of future hop uses are simply updated in the encapsulated route according to the delay caused by other bundles using the same contact.
  • If delaying the bundle causes its transmission at a time later than the corresponding contact’s end time, a new route is computed from the node where the problem happened, avoiding this problematic contact. The time of the new route-search call is the problem-detection time.
We provide further details regarding the second case: To reroute the bundle, the up-to-date CP at the problem detection time is loaded instantly from the source node. Note that this is an advantageous assumption for the benchmark. Moreover, since the remaining part of the old route is no longer used, the volumes of future contacts in the old route are updated locally in the CP at this node (i.e., incrementing the available volume of the unused contacts by the size of the bundle). CP updates performed on intermediate nodes are not reported to the source. The newly computed route, if any, is encapsulated within the bundle for further forwarding.
If none of the $k$ shortest routes is valid, the bundle is sent instantly to the source and rerouting from the source is performed after one time step.
V-2 Buffer management after the route search
During forwarding phase, we assume that a node only has access to the state of its buffer and that of its neighbors (nodes in contact). When a bundle is to be transmitted, the current node verifies if there is enough room in the buffer of the receiving node. If this is the case, the next node’s buffer occupancy is increased to book space for the bundle and the bundle is transmitted.
A waiting time is evaluated to determine how long the bundle will remain in the buffer of the next node. This time is equal to the transmission time from the current node to the next node $t_{\text{tx,1}}$, as well as the wait time until the start of the next contact (if this wait exists) $t_{\text{wait}}$ and the transmission time from the next node to a more distant node in the route $t_{\text{tx,2}}$. During $t_{\text{tx,1}}+t_{\text{wait}}$, the total bundle size is assumed to occupy the buffer of the next node. During $t_{\text{tx,2}}$, the space occupied by the bundle is linearly reduced as time passes (i.e., as the bundle leaves the buffer) according to the data rate of the contact used for this transmission.
If there is no room for the bundle in the buffer of the next node, the bundle is rerouted from the current node via a route that avoids this problematic neighbor as well as previously visited nodes to avoid loops. Note that a node whose buffer is fully occupied is considered problematic and must therefore be avoided. To do this, the problematic contact that was to be used in the route is removed from the CP for the rerouting of this bundle from this node only. Note that the bundle can pass through the problematic node later in the route, if the buffer has been emptied for example.
If no valid route is found among the $k$ shortest routes, the bundle is instantly sent back to the source and rerouted after one time step.
An alternative solution, which we are not considering, would be to continue trying to transmit via the fully booked node each time step until a valid route is found or a place becomes available in the buffer of the problematic node, whichever comes first. Note that this solution requires a lot of computing resources as the route search is repeated at each time step until a suitable solution is found.
VI Simulation environment and setup
VI-A Environment
[…]
We consider two GSs (GS 1 and GS 2). Bundles are generated by GS 1, which is the source. The bundles are to be transmitted from this source to GS 2, the destination. The satellites are the intermediate nodes of the network.
The CP is constructed for a 24-hour scenario starting from 8:55 pm on August 19, 2020. We assume that all the contacts have the same rate $R_{C_{j}}$.
A snapshot of the GSs and one Walker constellation with $N_{\text{sat}}=16$ at 11:18:08 pm on August 19, 2020 is illustrated in Figure 9. It was generated using MATLAB Satellite Communications Toolbox [22]. Using cislunar constellations to evaluate the algorithm will be considered in future work (see also the conclusions).
Figure 9: Snapshot of the satellite communication scenario with $N_{\text{sat}}=16$ at 11:18:08 pm on August 19, 2020. This corresponds to constellation 3 in Table I.
In the considered constellations, each orbital plane has $\frac{N_{\text{Sat}}}{N_{\text{p}}}$ satellites. We define the plane hops as the number of planes a communication between two satellites can cross (Table I). Depending on the communication parameters and the satellite dynamics, a satellite located in an orbital plane can communicate with satellites located in the same plane (zero plane hop, same RAAN), in adjacent planes (one plane hop, one $\Delta\Omega$ difference between RAANs of the satellites) or in more distant planes (more than one plane hops, meaning that there is no plane restriction in our case as the number of planes is limited).
The characteristics of the used constellations, also depending on the communication parameters provided in the following subsection, are given in Table I.
VI-B Communication parameters
Each satellite is mounted with transmit and a receive Gaussian antennas having a dish diameter of $0.5$ and an aperture efficiency of $0.65$. The transmission frequency and power are $30\times 10^{9}$ Hz and $15$ dB, respectively. The receiver’s gain to noise ratio and energy to noise sensitivity are equal to $3$ dB/Kelvin and $4$ dB, respectively. The source GS is equipped with a transmit antenna whose properties are similar to those of satellites, except that the transmission power is considered to be $30$ dB. The destination GS is equipped with a receive antenna whose properties are similar to those of the satellites, with the exception of its sensitivity of $1$ dB. A free space channel model is considered. The considered minimum required elevation angles of the GSs for communication is $0\degree$.
Two satellites can communicate with each other if they are in line-of-sight and if the power received by the receiving satellite from the transmitting satellite is greater than the sensitivity of the receiving satellite. For a satellite to communicate with a GS, the elevation angle of the former relative to the latter must be greater than the minimum elevation angle of the latter, in addition to being in line-of-sight and within the communication range. If these conditions are met, there is link closure between two nodes. A contact extends from the beginning to the end of a link closure between two nodes.
Example. In Figure 9, constellation $3$ is represented. The figure illustrates, in green, link closures between pairs of nodes: sat13-sat7, sat11-sat1, and sat14-GS 2, for example. These link closures happens at the time of the snapshot. The link between sat14 and GS 2 is at its limit: One time sample (of one second) later, this link no longer exists as these two nodes cease to respect the elevation angle constraint. The link closure is thus interrupted even though they remain in line-of-sight and in their communication range. In addition, sat4 and sat8 may be within communication range, but their communication is blocked by the Earth, preventing them from being in line-of-sight.
TABLE I: Constellation characteristics.
Constellation $N_{\text{p}}$ $N_{\text{sat}}$ Number of Mean contact Plane
number contacts in CP duration (s) hops
1 $4$ $4$ $60$ $511.03$ $1^{\text{(a)}}$
2 $4$ $8$ $175$ $667.33$ $2^{\text{(b)}}$
3 $4$ $16$ $767$ $778.36$ $2$
${}^{\text{(a)}}$ 1: A satellite can only communicate with satellites in same or
          adjacent orbital planes to its own plane (up to one plane hop).
${}^{\text{(b)}}$ 2: A satellite can also communicate with satellites in orbital planes
          adjacent to those adjacent to its own plane (up to two plane hops).
VI-C Other simulation parameters and assumptions
At each time step, one bundle of $100$ Bytes is generated at the source. The bundles are generated over a period of $500$ seconds, starting from the beginning of the scenario time, with equal inter-arrival times. In total $N_{\text{b}}$ bundles are generated and forwarded. This means that one bundle is generated each $\frac{500}{N_{\text{b}}}$ seconds (the bundle inter-arrival time decreases as $N_{\text{b}}$, the total number of bundles generated at source, increases).
Yen’s constant $k$ is chosen equal to $10$. Finally, we assume that there is no external disruption in the network. That is, the provided CP is valid and no rerouting is needed on intermediate nodes due to external factors. Given these assumptions, there is therefore no need to consider any safety margin for contact capacity and buffer space (Section IV-C).
VI-D Simulation tools
We use MATLAB Satellite Communications Toolbox [22] to simulate the presented satellite communication scenario. We can then calculate the link closures between each pair of nodes as a function of time to obtain the CP used in the benchmark and in our solution, which are also implemented in MATLAB.
An alternative algorithmic implementation would be to use Interplanetary Overlay Network (ION) [23] [24]. However, we have not chosen it because our main focus is on CGR.
VII Simulation results
We evaluate the performance using two main metrics. The first metric is the average time a bundle spends in the network. This is computed as the difference between the destination-arrival time and the bundle-generation time at the source. It includes buffer waiting time, transmission time, and potential rerouting delays. The other metric is the number of rerouting events. For SR-CGR, it includes rerouting at the source when no route is found among the $k$ shortest ones as well as when a bundle is instantly sent back to this source. It also includes, both for SR-CGR and the proposed solution, rerouting at an intermediate node if a buffer or a capacity check fails.
We evaluate the impact of contact-capacity management and buffer management separately. Of course, the two management techniques can be used together to obtain greater gain.
VII-A Impact of contact-capacity management
In this section, we only evaluate the performance of the proposed contact-capacity management. Therefore, we do not consider the constraints associated with buffer limits, i.e., the maximum node buffer capacity $B_{\text{max}}=~{}\infty$ for all nodes.
Figure 10 shows the average time spent in the network for $N_{\text{sat}}=16$ (constellation $3$) as function of the number of bundles generated at source $N_{\text{b}}$. Three values of contact data rate $R_{C_{j}}$ are considered: $200$, $400$, and $800$ bps for all contacts. As expected, one can see that the time spent in the network increases with the number of bundles and with decreasing data rate. With low load (e.g., $N_{\text{b}}<300$ for $400$ bps), the proposed algorithm and the SR-CGR have similar performance. As $N_{\text{b}}$ increases, the gap between the curves increases, with less time spent in the network using the proposed method. This is due to the increased number of rerouting events during forwarding time when using SR-CGR. This number of rerouting events is shown in Figure 11 for the same $N_{\text{sat}}$ (constellation $3$) and $R_{C_{j}}$ values as function of $N_{\text{b}}$. With low load, there are only few overbooking events happening on the contacts. In this case, traditional SR-CGR has satisfactory performance. When the number of bundles increases, the overbooking phenomenon increases, making rerouting more likely. In case of such overbooking event, intermediate nodes try to overcome this issue by rerouting the bundle (from their location toward the destination). Such new routes are longer than the ones avoiding overbooking (computed using the proposed algorithm). Moreover, if no route is found from the node, the bundle has to be rerouted from the source (see the assumptions on the benchmark) also causing additional delay. The proposed algorithm is more efficient because the shortest route is always computed directly from the source node, always avoiding overbooking. Indeed, as expected, no rerouting was required with our proposed solution in the simulation due to contact capacity check failures.
Figure 10: Average time spent in the network ($N_{\text{sat}}=16$).
Moreover, although in some cases the time spent in the network with SR-CGR is similar to the proposed solution, the node complexity is significantly higher with the benchmark due to multiple checks and rerouting events.
Figure 11: Number of reroute calls for CGR ($N_{\text{sat}}=16$).
Then, Figure 12 illustrates the time spent in the network with $4$, $8$, and $16$ satellites (Constellations $1$, $2$, and $3$, respectively). Contact data rate is considered equal to $400$ bps. With four satellites (one satellite per orbital plane), the number of nodes is limited and they are more distant compared to constellations with higher number of satellites. Thus, the number of communication opportunities are limited, limiting the routing degrees of freedom. In this case, the proposed algorithm does not show improvement compared to the benchmark. When the number of satellites increases to $8$ or $16$, our algorithm presents improvement compared to the benchmark, in particular with high load ($N_{\text{b}}>250$).
Figure 12: Average time spent in the network ($R_{C_{j}}=400$ bps).
VII-B Impact of buffer management
In this section, we evaluate the performance of the proposed buffer management. No contact capacity restriction is implemented, i.e., $R_{C_{j}}=\infty,\;\forall j$.
The average time spent by a bundle in the network as a function of the buffer size per node, expressed by the number of bundles a buffer can store, is given in Figure 13. We consider $N_{\text{sat}}=16$ (constellation $3$). The buffer storage size is identical for all intermediate nodes. The source node has infinite buffer size. The results are shown for $N_{\text{b}}=250$ and $N_{\text{b}}=100$ bundles generated during $500$ seconds. As expected, when the storage capacity of buffers increases, the time spent in the network is reduced. The gain is significant with the proposed solution, especially when the buffer sizes are limited. To reach the performance of infinite buffer capacity for $N_{\text{b}}=100$, our solution requires buffers that can store $10$ bundles, to be compared with $40$ bundles (almost half of $N_{\text{b}}$) to achieve this performance with the benchmark. The difference is even higher with $N_{\text{b}}=250$, where to reach the performance of infinite buffer the buffer sizes should be $40$ and $200$ bundles (small box in the figure) using our solution and the benchmark, respectively.
Figure 13: Average time spent in the network as function of the buffer size ($N_{\text{sat}}=16$).
Similarly to the capacity case, these differences are due to the number of rerouting events on intermediate nodes as shown in Figure 14, also for $N_{\text{sat}}=16$ (constellation $3$), when using SR-CGR. Moreover, as expected, no rerouting had to be performed with the proposed solution due to a buffer check failure.
Figure 14: Number of reroute calls for CGR ($N_{\text{sat}}=16$).
Finally, Figure 15 shows the impact of different constellations, namely $1$ and $3$, on the performance related to buffer constraints and for $N_{\text{b}}=250$. When the number of nodes is limited, each of them is more frequently used. The limited storage capacity then induces additional delay for both solutions: With $N_{\text{sat}}=~{}4$, the two solutions reach the infinite buffer performance at the same buffer capacity ($100$ bundles per node). Nevertheless, with that same constellation and with lower buffer size per node, our proposed solution manages to significantly reduce the average time a bundle spends in the network. Increasing the number of nodes in the constellation and using our proposed solution leads to very low delay even with low buffer capacity per node.
Figure 15: Average time spent in the network as function of buffer size for $N_{\text{b}}=250$.
VIII Conclusions
In this work, we proposed enhanced contact-capacity management and buffer management to improve the state-of-the-art CGR routing algorithm. We proposed managing these resources by taking actions before or during the route search. Our solution is a global solution: It considers and tracks the evolution of the contact capacity and buffer occupation in all contacts and nodes of the network. The resulting route is optimal and respects contact capacity and node buffer constraints.
To manage contact capacity, we introduced contact splitting. It allows for consideration of the previously consumed portions of a contact by routed bundles. For the buffer management, we proposed temporary bundle-specific contact splitting and edge pruning to avoid any portion or succession of contacts that might result in buffer overflows. These temporary modifications depend on the forecast buffer tables that track the future states of the node buffer occupancy level as function of time.
We compared the performance of the proposed solution with the benchmark SR-CGR. The solution offers considerable gains in terms of time spent in the network and intermediate node complexity, measured as the number of rerouting calls.
Limitations. Although the proposed solution offers improvements over the state-of-the-art while meeting capacity and buffer constraints, it has some limitations that could be taken into account for further improvements. Our solution is based on SR-CGR, which is usually used to reduce CGR complexity. This reduction is achieved by performing one route search per bundle rather than multiple route searches (per hop routing). Nevertheless, our solution splits a contact into up to two contacts, when a bundle occupies a contact capacity or a buffer space. The splitting action increases the size of the CP, which directly impacts the complexity of the route search. In our solution, the additional complexity is tractable because as the route search is performed only on source nodes and not on all low-power intermediate nodes. Nevertheless, when moving to a distributed solution, one should consider such additional complexity. The centralized solution is also a limitation of this study. The examination of an efficient distributed solution is therefore an interesting research direction as improvement of this work.
Future work. We plan to assess proposed cislunar constellations, from a communication perspective, using the presented routing technology. We believe that the proposed routing algorithm may help the design of the communication and storage systems to be embedded in the satellites of such lunar constellations.
We also plan to investigate traffic-level solutions rather than a per-bundle solution. In the state-of-the-art as in this work, the route is chosen based on optimal criterion at the bundle level, not for the entire traffic to be transmitted. Therefore, this per-bundle strategy is likely to be not optimal for the overall communication performance.
Finally, as mentioned in the previous “limitations” subsection, investigating distributed solutions rather than a centralized one is also of interest.